OpenStudy (agar.io):

(4/5)x = 2/5 @TheSmartOne I got x = 1/2

OpenStudy (anonymous):

you're right.

TheSmartOne (thesmartone):

@Agar.io @~ You are correct! :) \[\frac{5}{4}\times \frac{2}{5} = \frac{2}{4} = \boxed{\frac{1}{2}}\] So, correct! :)